drtyguhi7y86t5rrftyu:

PLEASE HELP!!!!!!!! A candy bar box is in the shape of a triangular prism. The volume of the box is 1,200 cubic centimeters. A triangular prism is shown with base of triangle labeled 10 cm, sides of the triangles labeled 13 cm, and length of the box equal to 20 cm. Part A: What is the height of the base? Show your work. (5 points) Part B: What is the approximate amount of cardboard used to make the candy box? Explain how you got your answer.

drtyguhi7y86t5rrftyu:

TETSXPREME:

60 Area260

akeelahmeans:

Using the Pythagorean Theorem we split the base in 2 which gives us 5 and we start the process 5^2+x^2=13^2 25+x^2=169 -25+x^2=169-25 x^2=144 take the square root of both sides and we get 12 as the height. Doing b*h/2 to get the area of the triangular sides we get 12*10=120/2 we get 60 as the amount of cardboard they need to use. For the other side, we use l*w=A to get 260. Hope this helps! :D

drtyguhi7y86t5rrftyu:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @akeelahmeans Using the Pythagorean Theorem we split the base in 2 which gives us 5 and we start the process 5^2+x^2=13^2 25+x^2=169 -25+x^2=169-25 x^2=144 take the square root of both sides and we get 12 as the height. Doing b*h/2 to get the area of the triangular sides we get 12*10=120/2 we get 60 as the amount of cardboard they need to use. For the other side, we use l*w=A to get 260. Hope this helps! :D \(\color{#0cbb34}{\text{End of Quote}}\) ty

akeelahmeans:

No problem